# appReferences in the Appendix

### v2

@@ 3.3 Algorithm Extension
Building on Algorithm 3, we propose an enhanced multi-stage learning algorithm that employs the first-order method for decision-making during each re-solving interval. This multi-restart mechanism further improves the accuracy and efficiency of our framework. We illustrate this algorithm in the following figure.
Figure 2: Enhanced Parallel Multi-Phase Online Learning and Decision-Making Algorithm
Figure 2 illustrates this variant of Algorithm 3 with multi-stage exploration and exploitation. Rather than make decisions solely based on the most recently solved dual prices, the algorithm treats the updated dual prices at each re-solving interval as a new starting point for the first-order learning process. For example, in Algorithm 3, the dual price $\bm{p}_{f}^{L}$ is used to guide decision-making during the time interval $t\in[f,2f]$. In contrast, the enhanced algorithm restarts the first-order method by setting $\bm{p}_{f}^{L}=\bm{p}_{f}^{D}$ and then continues to fine-tune the dual prices $\{\bm{p}_{t}^{D}\}_{t=f}^{2f}$ to make up-to-date decisions. Algorithm 4 summarizes this approach.
Algorithm 4 Enhanced Parallel Multi-Phase Online Learning and Decision-Making
Input: total resource $\bm{b}$, time horizon $T$, average resource $\bm{d}=\bm{b}/T$, initial dual price $\bm{p}_{1}=\bm{0}$, re-solving frequency $f$, and number of re-solving batches $k=\lfloor T/f\rfloor$.
for $t=$ $1$ to $T$  do
       Observe $(r_{t},\bm{a}_{t})$ and make decision $x_{t}$ based on rule (6). Update constraint for $i=1,...,m$:       remaining resource constraint $b_{it}=b_{i,t-1}-a_{it}x_{t}$       average resource capacity $d_{it}=\frac{b_{it}}{T-t}$ Update learning rate $\alpha_{t}=\mathcal{O}(1/t)$. if $t\bmod f\neq 0$ then
             Compute the subgradient and update dual price $\bm{p}_{t+1}$:
$\bm{p}_{t+1}$ $=(\bm{p}_{t}-\alpha_{t}(\bm{d}-\bm{a}_{t}x_{t}))\vee\bm{0}$
@@
             Solve the online dual problem and update dual price $\bm{p}_{t+1}$:
$\bm{p}_{t+1}=\operatornamewithlimits{arg\,min}_{\bm{p}\geq\bm{0}}\ \bm{d}_{t}^{\top}\bm{p}+\frac{1}{t}\sum_{j=1}^{t}(r_{j}-\bm{a}_{j}^{\top}\bm{p})^{+}$
@@
This algorithm establishes an online framework of multi-stage learning and decision-making. The incorporation of an intermediate first-order method improves the algorithm’s stability and ensures smooth resource consumption. These improvements result in better performance during the final batch, thereby reducing the algorithm’s total regret.
@@ 3.4 Algorithm Application
Our motivation in designing Algorithm 3 is to develop a hybrid decision-making system that effectively balances computational efficiency and decision accuracy. By integrating the first-order method, we aim to reduce computational costs, while the LP-based method enhances the precision of our decisions. Building on the Spectrum Theorem 3.2, this section aims to translate our theoretical results into practical applications. Specifically, we formulate a new optimization problem to determine the optimal re-solving frequency, taking into account computational resource constraints.
We use $c_{1}(\cdot),c_{2}(\cdot)$ to denote the computational cost functions for the LP-based method and first-order method, respectively, and use $R$ to denote the total computational resource capacity. Then we define the following optimization problem to determine the optimal re-solving frequency $f$:
$\min_{f\in\mathbb{Z}^{+}}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \log\Big{(}\frac{T}{f}\Big{)}+\sqrt{f}$ (24)
subject to $\leavevmode\nobreak\ \quad c_{1}(k)+2c_{2}(f)\leq R$
$\leavevmode\nobreak\ \quad 1\leq f\leq T$
We can further analyze the computational requirements of our algorithm by considering the specific LP solvers employed. Recall that $m$ denotes the dimension of the dual price vector. At any given time step $t$, utilizing either the interior-point method or the simplex method as the LP-based solver incurs a computational cost of $m^{2}(m+t)$. In contrast, the first-order method incurs a computational cost of $2mf$. We encapsulate these trade-offs in the following proposition.
@@
Given a fixed computational resource capacity $R$, if we use the interior-point method or simplex method as the LP solver in Algorithm 3, we can formulate the optimization problem (24) as
$\min_{f\in\mathbb{Z}^{+}}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \log\Big{(}\frac{T}{f}\Big{)}+\sqrt{f}$ (25)
subject to $\leavevmode\nobreak\ \quad\sum_{b=1}^{k}\left(m^{2}(m+bf)\right)+2mf\leq R,$
$\leavevmode\nobreak\ \quad 1\leq f\leq T$ (26)
By solving optimization problem (25), we can select an optimal re-solving frequency $f$ that achieves the lowest regret within the computational resource limit $R$.
Therefore, we extend our original optimization problem for resource allocation to include computational resource allocation. This proposition enables users to determine the optimal re-solving frequency that balances regret with computational cost based on their available computational resources.
@@ 4 Numerical Experiments
We conduct extensive experiments under various settings to illustrate the performance of our algorithm and validate the theoretical results of our framework. This section is divided into two parts. In the first part (Section 4.1), we evaluate Algorithms 3 and 4 across different choices of the re-solving frequency. In the second part (Section 4.2), we compare our algorithm with the LP-based method in terms of running time and with the first-order method in terms of regret. Both sets of experiments are conducted under different distributions of $\{(r_{t},\bm{a}_{t})\}_{t=1}^{T}$ that satisfy the assumptions in Section 2.
4.1 Performance Evaluation under Varying Re-solving Frequencies
We generate the sequence $\{(r_{t},\bm{a}_{t})\}_{t=1}^{T}$ randomly from distributions that satisfy Assumptions 1 and 2. Specifically, we use the uniform and normal distribution, with details in Table 2. We select the types of resource $m\in\{1,5\}$, time horizon $T$ evenly spanned over $[10^{2},10^{6}]$, and initial average resource capacity $d_{i}$ sampled i.i.d. from $\text{Uniform}[1/3,2/3]$. We run our algorithm for $100$ trials and use the average as the final result for each experiment.
Table 2: Stochastic Input in Experiments
Distribution Model $\bm{a}_{t}$ $r_{t}$
Input I $a_{it}\sim\text{Uniform}[0,2]$ $r_{t}\bot\bm{a}_{t},r_{t}\sim\text{Uniform}[0,10]$
Input II $a_{it}\sim\text{Normal}(0.5,1)$ $r_{t}=\sum_{i=1}^{m}a_{it},r_{t}\sim\text{Normal}(0.5m,m)$
We evaluate our algorithm with $m=1$ under the stochastic input models I and II from Table 2. The experiments are conducted across various choices of $f\in\{T^{1/3},T^{1/2},T^{2/3}\}$ representing high, medium, and low re-solving frequencies. We analyze the absolute value of the regret and its growth rate over time. We illustrate our results in Table 3 and Figure 3 on a logarithmic scale.
Table 3: Regret of Algorithms with various re-solving frequency
Offline Low freq Mid freq High freq
$T=10^{3}$ 12.13 7.76 5.77 4.86
$T=10^{4}$ 38.50 10.96 7.55 5.67
$T=10^{5}$ 122.44 23.90 9.92 8.36
$T=10^{6}$ 404.59 56.70 21.90 8.99
(a) Algorithm 3, Input I
Offline Low freq Mid freq High freq
$T=10^{3}$ 11.44 6.28 4.86 3.95
$T=10^{4}$ 36.50 10.21 7.34 3.81
$T=10^{5}$ 115.57 14.61 11.78 4.66
$T=10^{6}$ 365.99 35.20 15.68 6.26
(b) Algorithm 3, Input II
Offline Low freq Mid freq High freq
$T=10^{3}$ 12.13 6.78 5.20 4.50
$T=10^{4}$ 38.50 10.37 8.03 5.99
$T=10^{5}$ 122.444 22.33 11.57 6.36
$T=10^{6}$ 404.59 48.21 22.44 7.09
(c) Algorithm 4, Input I
Offline Low freq Mid freq High freq
$T=10^{3}$ 11.44 3.20 2.56 1.75
$T=10^{4}$ 36.50 5.48 4.30 2.52
$T=10^{5}$ 115.57 12.35 4.48 3.86
$T=10^{6}$ 365.99 30.48 13.20 4.77
(d) Algorithm 4, Input II
(a) Algorithm 3 with input I
(b) Algorithm 3 with input II
(c) Algorithm 4 with input I
(d) Algorithm 4 with input II
Figure 3: Main Algorithms with Various Frequency
These experiments assess the performance of our two main algorithms under varying re-solving frequencies and across multiple time horizons. As presented in Table 3 and Figure 3, we observe that regret decreases as the re-solving frequency increases from $f=T^{2/3}$ to $f=T^{1/3}$. This trend holds consistently across both algorithms and input types, with higher re-solving frequencies achieving the lowest regret in all cases. In addition, while regret accumulates over longer time horizons, the rate of increase remains stable for algorithms employing higher re-solving frequencies. These findings illustrate the robustness of our algorithms, as more frequent updates enable better adaptation to dynamic environment.
4.2 Comparative Analysis with Baseline Methods
After examining the algorithm under various re-solving frequencies, we compare its performance with the classic first-order method and LP-based method in terms of regret and computation time. We generate $\{(r_{t},\bm{a}_{t}\}_{t=1}^{T}$ from the uniform distribution under the Input I model (Table 2). We set the number of resource types to $m=5$, the time horizon $T$ to range evenly over $[10^{2},10^{6}]$, and the average resource capacity $d_{i}$ to be sampled i.i.d from $\text{Uniform}[1/3,2/3]$. The re-solving frequency is fixed to be $f=T^{1/3}$. Each result is averaged over 100 trial runs, and the findings are summarized in Table 5 and Figure 5.
@@ 4.2 Comparative Analysis with Baseline Methods
$10^{3}$ $66.88$ Offline $0.001$
$2.80$ LP-based (Algorithm 1) $4.185$
$25.11$ First-Order (Algorithm 2) $0.002$
7.20 Algorithm 3 0.3
4.50 Algorithm 4 0.3
$10^{4}$ 115.32 Offline $0.008$
3.50 LP-based (Algorithm 1) $123.497$
60.39 First-Order (Algorithm 2) $0.013$
@@
$10^{5}$ 203.20 Offline $0.118$
3.88 LP-based (Algorithm 1) $>3600$
73.07 First-Order (Algorithm 2) $0.108$
@@ 3.88 LP-based (Algorithm 1) $>3600$
$10^{6}$ 351.91 Offline 1.211
5.50 LP-based (Algorithm 1) $>100000$
115.39 First-Order (Algorithm 2) 1.577
@@ 5.50 LP-based (Algorithm 1) $>100000$
Table 4: Algorithms Comparison.
Figure 5: Regret for Various Algorithms
Table 5 and Figure 5 provide a comprehensive comparison of all the algorithms discussed in this paper. Focusing on our primary algorithms, Algorithm 3 and 4, we observe:
@@
    Our algorithms exhibit strong performance in terms of decision optimality. They achieve over a 20-fold improvement in regret compared to the offline solution and more than a $10$-fold improvement compared to the first-order method. These numerical results also corroborate the theoretical bounds in Theorem 3.2.
@@
    Our algorithms are computationally efficient. Their computation times exhibit a 100-fold improvement compared to the LP-based method while maintaining regret levels that are remarkably close to those of the LP-based method. This significant improvement in efficiency makes our algorithms highly suitable for real-time and large-scale applications.
Therefore, we achieve a balance between effective decision-making and efficient computation. Our algorithms demonstrate better regret than the first-order method and obtain substantial computational speed-ups compared to the LP-based method. Our algorithms exhibit robust scalability and adaptability under various re-solving frequencies and stochastic input models, consistently maintaining superior performance as the problem size grows. This adaptability ensures their effectiveness in diverse scenarios.
5 Conclusion
This paper addresses an online linear programming problem for decision-making in resource allocation. Identifying existing methods’ limitations, we develop a novel two-path framework that decouples online learning and decision-making into independent processes. Our cohesive framework integrates the LP-based method and first-order method, thereby balancing decision-making optimality and computational efficiency. We prove that our algorithm attains $\mathcal{O}(\log(T/f)+\sqrt{f})$ worst-case regret under continuous distributions, and our experimental results demonstrate strong regret and runtime improvements over competitive baselines.

### v3

@@ 3.3 Algorithm Extension
Figure 2: Algorithm 2 illustration of parallel paths with multi-time restart. Decisions are generated based on 1) the LP-based method (blue) with frequency $f$ and 2) the first-order method (red) during re-solving intervals with a warm start.
We propose an enhanced version of Algorithm 1 that employs the first-order method for decision-making between two consecutive LP resolves. Rather than making decisions solely with the most recently solved dual price from the learning path, Algorithm 2 treats it as a new starting point for each re-solving interval and adopts a smaller step size for subgradient updates to avoid deviating too far from the LP-guided solutions. Figure 2 illustrates this framework. Specifically, for each interval $[jf,(j+1)f]$, we attain the new start $\bm{p}_{jf}$ from the learning path at time $jf$ and continue to fine-tune the dual price $\{\bm{p}_{t}\}_{t=jf}^{(j+1)f}$ using the first-order method with a learning rate $\alpha_{t}=\mathcal{O}(1/t)$. Algorithm 2 summarizes this approach.
As we illustrate in our experiments (Section 4), the incorporation of an intermediate first-order method improves the algorithm’s stability and ensures smooth resource consumption. Therefore, the multi-restart mechanism results in better performance during the final batch and improves the algorithm’s total performance.
Algorithm 2 Enhanced Multi-Start OLP Algorithm
Input: total resource $\bm{b}$, horizon $T$, average resource $\bm{d}=\bm{b}/T$, initial dual $\bm{p}_{1}=\bm{0}$, re-solving frequency $f$
for $t=$ 11 to $T$ do
   Observe $(r_{t},\bm{a}_{t})$ and make decision $x_{t}$ as rule (6) Update constraint for i=1,,mi=1,...,m:   remaining resource constraint $b_{it}=b_{i,t-1}-a_{it}x_{t}$   average resource capacity $d_{it}=\frac{b_{it}}{T-t}$ if $t\bmod f\neq 0$ then
      Update learning rate $\alpha_{t}=\mathcal{O}(1/t)$ Compute subgradient and update dual price $\bm{p}_{t+1}$:
$\bm{p}_{t+1}$ $=(\bm{p}_{t}-\alpha_{t}(\bm{d}-\bm{a}_{t}x_{t}))^{+}$
@@
      Solve OLP and update dual price $\bm{p}_{t+1}$:
$\bm{p}_{t+1}=\operatorname*{argmin}_{\bm{p}\geq\bm{0}}\ \bm{d}_{t}^{\top}\bm{p}+\tfrac{1}{t}\textstyle\sum_{j=1}^{t}(r_{j}-\bm{a}_{j}^{\top}\bm{p})^{+}$
@@ 3.4 Algorithm Application
Our motivation in designing Algorithm 1 is to effectively balance computational efficiency and decision optimality. Building on the Spectrum Theorem 3.2, this section aims to translate our theoretical results into practical applications.
We formulate a new optimization problem to find the optimal re-solving frequency that minimizes the regret in ˜3.2, taking into account computational resource capacities. Denote $c_{1}(\cdot),c_{2}(\cdot)$ as computational cost functions for LP-based and first-order methods, and $R$ as the total computational resource capacity. The optimal value of $f$ is determined by solving the following optimization problem:
$\min_{f\in\{1,\dots,T\}}$ $\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \log\Big(\tfrac{T}{f}\Big)+\sqrt{f}$
s.t. $\penalty 10000\ \quad c_{1}(k)+2c_{2}(f)\leq R.$ (14)
Specifically, if we use the interior-point method or the simplex method as the LP solver, the computational cost is $m^{2}(m+t)$ for any time $t$. The first-order method updates gradients in constant time. The following proposition provides a concrete example in practice.
@@
Given a fixed computation resource capacity $R$, if we use the interior-point or simplex method as the LP solver in Algorithm 1, we can instantiate Constraint (14) as
$\textstyle\sum_{j=1}^{k}(m^{2}(m+jf))+2mf\leq R.$
This proposition enables users to determine the optimal re-solving frequency that balances regret and computational cost based on available computational resources.
@@ 4 Numerical Experiments
Figure 3: Algorithms comparison.
@@
$0^{4} 11$5.32 $\mathcal{O}(T^{1/2})$ First-Order $0.008$
60.39 $\mathcal{O}(T^{1/3})$ First-Order $0.013$
3.50 LP-based $123.497$
@@
$10^{5}$ 203.20 $\mathcal{O}(T^{1/2})$ First-Order $0.118$
73.07 $\mathcal{O}(T^{1/3})$ First-Order $0.108$
3.88 LP-based >3600>3600
@@
$10^{6}$ 351.91 $\mathcal{O}(T^{1/2})$ First-Order $1.211$
115.39 $\mathcal{O}(T^{1/3})$ First-Order $1.577$
5.50 LP-based >100000>100000
@@
Figure 5: Regret for various algorithms.
We conduct extensive experiments to evaluate our algorithm’s performance and validate our theoretical results. This section is divided into two parts. In the first part (Section 4.1), we evaluate Algorithms 1 and 2 across different choices of re-solving frequency. In the second part (Section 4.2), we compare our algorithm with LP-based and first-order methods in terms of regret and running time. All implementations can be found at GitHub Link.
We consider the following distributions:
$\text{Input I:}\ \ a_{it}\sim\text{Unif}[0,2],r_{t}\sim\text{Unif}[0,10]$
$\text{Input II:}\ \ a_{it}\sim\mathcal{N}(0.5,1),r_{t}\sim\mathcal{N}(0.5m,m)$
Learning rates are selected as specified in Section 3:
$\text{Algorithm 1: }\alpha_{t}=\begin{cases}\mathcal{O}(1/f^{1/2})&t\leq f\\ \mathcal{O}(1/f^{2/3})&t\geq kf\end{cases}$
$\text{Algorithm 2: }\alpha_{t}=\mathcal{O}(1/t)$
4.1 Regret under Varying Re-solving Frequencies
We choose $m=1$ and generate the sequence $\{(r_{t},\bm{a}_{t})\}_{t=1}^{T}$ randomly from the uniform (Input I) and normal (Input II) distributions, and include more complex distributions in Appendix E.1. The time horizon $T$ spans evenly over $10^{2},10^{6}]$, the initial average resource is sampled as $d_{i}\sim\text{Uniform}[1/3,2/3]$, and $f\in\{T^{1/3},T^{1/2},T^{2/3}\}$ representing high, medium, and low re-solving frequencies. We report the average result over 100 trials for each experiment. We use the classic first-order method with $\mathcal{O}(T^{1/2})$ regret (Li et al., 2020) as a baseline. Results are summarized in Table 1 and Figure 4 on a logarithmic scale.
We analyze the absolute value of regret and its growth rate over time. We observe that regret decreases as the re-solving frequency increases. This trend holds consistently across both algorithms and input types. In addition, while regret accumulates over longer time horizons, the increasing rate remains stable for algorithms employing higher re-solving frequencies. These findings are consistent with the guarantees of ˜3.2, as more frequent updates enable better adaptation to dynamic environments.
4.2 Comparative Analysis with Baseline Methods
Table 1: Regret of algorithms with various re-solving frequencies.
$T$ First-Order Low freq Mid freq High freq
Input I $10^{3}$ 12.13 7.76 5.77 4.86
$10^{4}$ 38.50 10.96 7.55 5.67
$10^{5}$ 122.44 23.90 9.92 8.36
$10^{6}$ 404.59 56.70 21.90 8.99
Input II $0^{3} 11$.44 6.28 4.86 3.95
$10^{4}$ 36.50 10.21 7.34 3.81
$0^{5} 11$5.57 14.61 11.78 4.66
$10^{6}$ 365.99 35.20 15.68 6.26
(a) Algorithm 1
$T$ First-Order Low freq Mid freq High freq
Input I $10^{3}$ 12.13 6.78 5.20 4.50
$10^{4}$ 38.50 10.37 8.03 5.99
$10^{5}$ 122.44 22.33 11.57 6.36
$10^{6}$ 404.59 48.21 22.44 7.09
Input II $0^{3} 11$.44 3.20 2.56 1.75
$10^{4}$ 36.50 5.48 4.30 2.52
$0^{5} 11$5.57 12.35 4.48 3.86
$10^{6}$ 365.99 30.48 13.20 4.77
(b) Algorithm 2
Figure 4: Evaluations of Algorithm˜1 and 2 across various horizons, re-solving frequencies, and stochastic inputs, validating the positive relationship between regret and frequency stated in ˜3.2.
We next compare our algorithm’s regret and computation time with a classic LP-based (Li & Ye, 2022) and two first-order methods with $\mathcal{O}(T^{1/2})$ (Li et al., 2020) and $\mathcal{O}(T^{1/3})$ regrets (Gao et al., 2024). We generate $\{(r_{t},\bm{a}_{t})\}_{t=1}^{T}$ from a uniform distribution (Input I). We set the resource types to $m=5$, time $T$ to range evenly over $10^{2},10^{6}]$, average resource $d_{i}\sim\text{Uniform}[1/3,2/3]$, and re-solving frequency to $f=T^{1/3}$ from Section 4.1. Each result is averaged over 100 trial runs. We summarize the findings in Table 5 and Figure 5, and we include more comparisons with an infrequent re-solving method (Li et al., 2024) in Appendix E.2. We observe the effectiveness of Algorithm˜1 and 2:
@@ 4.2 Comparative Analysis with Baseline Methods
    Our algorithms exhibit strong performance in terms of decision optimality. They achieve over a 20-fold and 10-fold improvement in regret compared to the $\mathcal{O}(T^{1/2})$ first-order method and the $\mathcal{O}(T^{1/3})$ first-order method respectively. These numerical results also corroborate the theoretical bounds in Theorem 3.2.
@@
    Our algorithms are computationally efficient. Their runtimes exhibit over 100-fold improvement compared to the LP-based method with only a minimal increase (less than 2-fold) in regret.
Therefore, we achieve a balance between effective decision-making and efficient computation. Our algorithms demonstrate better regret than the first-order method and obtain substantial computational speed-ups compared to the LP-based method. They exhibit strong scalability and adaptability across various re-solving frequencies and stochastic input models, consistently delivering superior performance as the problem size grows.
5 Conclusion and Discussion
This paper presents a new approach to online linear programming for dynamic resource allocation, addressing the inherent trade-offs between decision-making optimality and computational efficiency. Recognizing the limitations of existing methods, we propose a parallel framework that decouples online learning and decision-making into independent yet complementary processes. By integrating LP-based and first-order methods, our framework effectively balances total regret and computational cost.
We establish rigorous theoretical guarantees, proving that our algorithm achieves a worst-case regret bound of $\mathcal{O}(\log(T/f)+\sqrt{f})$ under continuous support. This result highlights our method’s ability to interpolate between LP-based and first-order methods based on computational capability. Furthermore, extensive experiments validate the effectiveness of our approach, demonstrating improvements in both regret minimization and runtime efficiency over competitive baselines.
Beyond these contributions, our work opens avenues for further research in adaptive online decision-making. Future directions include refining the re-solving frequency based on real-time computational constraints and extending our framework to broader classes of online optimization problems. Overall, our results underscore the potential of hybrid algorithms in high-dimensional, large-scale environments, offering practical insights for applications in operations research, machine learning, and beyond.
Discussion of Unknown Horizon
In this paper, we consider decision-making under a finite horizon, since the average resource ˜2.1(c) widely used in OLP literature relies on a fixed horizon to be well-defined. When the horizon is unknown, this assumption becomes ill-posed, and prior work (Balseiro et al., 2023) shows that it may not be possible to achieve sublinear regret. Addressing these challenges would likely require first adapting LP-based and first-order OLP methods to uncertain horizons before tackling a hybrid approach.
In practice, it is often possible to make a prior prediction of the horizon using data-driven approaches or based on the resources available. For example, in online advertising, the number of customers can be predicted from historical statistics. Besides, in retail or event-based sales, the selling horizon may be externally decided by upper-level decision-makers. These practical applications suggest a reasonable modeling choice for finite time horizons.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Acknowledgments
This work was supported in part by NSF grant CCF-2338226.
@@ Acknowledgments
Structure of the Appendix
We organize the appendix as follows. In Section A, we present some primary results for the OLP problem and the foundation analysis for LP-based and first-order methods; Section B characterizes the unique properties from our parallel multi-phase structure, and presents the cohesive mathematical framework which decomposes the performance metrics to unify all the OLP methods; Section C demonstrates the outline of Algorithm˜2, combines all the previous results and proves the main results in our paper; Section D provides auxiliary results and technical support for the previous three sections; Section E includes supplementary experiments of more general input distributions and comparisons with recent algorithms.
Appendix A Primary Results and Properties
In this section, we present some primary results for the problem and OLP dual algorithms which will help present our main results. We also include basic properties and convergence of the LP-based and first-order methods. All the analyses are stated under Assumption 2.1 and 2.2.
A.1 Preliminary results
We propose the following lemma as directed results from Assumptions. Let $\Xi$ denote the family of distributions satisfying ˜2.1 and 2.2. We propose the following lemma.
Lemma A.1 (Dimension Stability).
Under Assumptions 2.1 and 2.2, there exists a constant $\delta>0$ such that